Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
carbognani, gherardini, sposito, taranta - 18:59 Friday 01 July 2022 (56328) Print this report
Comment to Operator Report - Morning shift (56326)

As far as the problems reported to move the NI_HM_tx/ty picomotors we tryied:

  • stop/restart the TCSPicoNI server: no effects
  • reset (OFF/ON) the controller: no effects
  • swap WI and NI controllers: no effects

From the software/config point of view:

  • the ActuatorApp corresponding software (version v5r1p1) is unchanged from 2018
  • TCSPicoNI server config file is unchanged from 2016

We tryied to send command from a terminal via cm send just to exclude any problem from the Python client side but we had the same result.

The behaviour is that while moving both tx and ty channels in Forward direction, logs from the server report the move is performed but from the HWS we see no move at all,  while moving in Backward direction the move is reported as not performed by the controller. Correspondig logs below.

2022-07-01-12h23m11-UTC>INFO...-MoveRel NI_HM_ty F 5500.000000 command received from position -53337.000000
2022-07-01-12h23m15-UTC>INFO...-Memory Used increase, cur. 3884.00(KB), inc. 24.00(KB)
2022-07-01-12h23m19-UTC>INFO...-Ping command received
2022-07-01-12h23m40-UTC>INFO...-Ping command received
2022-07-01-12h23m51-UTC>INFO...-Ping command received
2022-07-01-12h24m09-UTC>INFO...-main> Stopped at position = -47837.000000

2022-07-01-10h30m06-UTC>INFO...-MoveRel NI_HM_tx B 5500.000000 command received from position -50365.000000
2022-07-01-10h30m08-UTC>INFO...-main> Stopped at position = -50365.000000

Check to hear the picomotor  noise from outside the TCS bench enclousure were inconclusive. Inspection on the bench itself  seems needed., 

Search Help